At the beginning of the month, you owned $8,000 of News Corp, $8,300 of First Data, and $8,700 of Whirlpool. The monthly returns for News Corp, First Data, and Whirlpool were 9.40 percent, −2.73 percent, and 11.53 percent. What’s your portfolio return? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Total value=(8000+8300+8700)=25000

Hence portfolio return=Respective return*Respective investment weight

=(8000/25000*9.4)+(8300/25000*-2.73)+(8700/25000*11.53)

=6.11%(Approx).